OpenStudy (anonymous):

Given the parallelogram EFGH, what is the length of side line EF. ( line EH - 3x+6) (line HG - 2x+5) ( line FG - 27) OPTIONS - A. 27 ; B. 21 ; C. 19 ; D. 7

OpenStudy (anonymous):

opposite sides of paralllelogram are equal so 3x+6 = 27 giving x = 7 EF = HG = 2x+5 so EF = 2(7) + 5 = 19 answer C - 19

OpenStudy (anonymous):

thanks